At the movie theatre, child admission is 5.30 an adult admission is 9.10. On Monday, 159 tickets were sold for a total sales of 1120.10, how many adult tickets were sold that day?
Answer by dabanfield(803)

At the movie theatre, child admission is 5.30 an adult admission is 9.10. On Monday, 159 tickets were sold for a total sales of 1120.10, how many adult tickets were sold that day?
Let a be the number of adult tickets and c the number of children's tickets. Then we have:
1.) a + c = 159
2.) 9.10*a + 5.30*c = 1120.10
From 1.) we have a = 159 - c.
Substitute 159 - c for a in equation 2.)and solve for c:
9.10*(159-c) + 5.30*c = 1120.10
Once the value for c is determined then a = 159 - c.